# Large Language Models for Relevance Judgment in Product Search

### v1

@@ 1. Introduction
Our results in section (3) show that Low Rank Adaptation (LoRA) (Hu et al., 2022) of intermediate-sized models such as Llama2 7B and Mistral 7B with the training datasets sized in the order of $10^{6}$ perform better than full model finetuning. This observation emphasizes the tradeoff between the model and data size. For smaller training datasets, LoRA is expected to align with data more efficiently than full finetuning. We also find that including additional item textual data in the form of description, as expected, improves performance in larger multi-billion parameter models, but not in smaller ones of size in the order of hundred million parameters.
Our training dataset is comprised of $6,055,251$ QIPs, $708,895$ queries and $2,664,741$ items. The dataset, similar to those in the public domain (e.g. (Reddy et al., 2022)), is imbalanced towards the most relevant label, but unlike the ESCI dataset introduced in (Reddy et al., 2022)) has an almost complete coverage of longer item description information. [^1] The dataset introduced in (Reddy et al., 2022) has an “item_description” coverage close to 48%. Our relevance models are classifiers that are trained on inputs of concatenated item textual features and query along with relevance labels. On a three-class graduation of relevance labels (relevant, related, irrelevant) we achieve micro f1 scores of 88.8% with Llama2 7B and 89.6% with Mistral 7B models. There are improvements up to 4.6% in f1 micro scores compared to the performance of pretrained off-the-shelf decoder-only LLMs (see tables 6, 7, and 8).
@@
We also demonstrate the regions of LoRA rank and $\alpha$, as well as LoRA dropouts, where Llama2 and Mistral 7B perform best, showing that, unlike conventional wisdom, with the size of the training dataset and LLMs in question, $\alpha$s in the range of 1 to 2 times LoRA rank perform better than alternatives. LoRA finetuning, in turn, results in higher f1 scores than full model finetuning in our trainings.
@@ 3.2. Relevance Models
We have employed 7B versions of Llama (Touvron et al., 2023) and Mistral (Jiang et al., 2023) in our training along with BERT/DeBERTa (Devlin et al., 2018) (Abolghasemi et al., 2022), (Reimers and Gurevych, 2019),(He et al., 2023a) as a baseline for larger models such as Llama and Mistral.
@@
In the following, we present the model training procedure for BERT/DeBERTa, Llama and Mistral.
3.2.1. BERT
@@
Previous studies have shown that the cross encoder based BERT model (Luan et al., 2021) performs well on solving semantic retrieval problems. To address the specific challenge of e-commerce product relevance search, we have employed a cross encoder model in our design. We experimented with finetuning our model using two variations of the BERT model as the starting checkpoints: 1) the BERT-Base uncased model from Huggingface, and 2) this same model further pretrained on the Walmart Search corpus.
@@
Walmart Pretrained BERT: Since the BERT-base model was trained on general domain text data, we aimed to enhance its performance on e-commerce related texts by conducting additional language model training using Walmart product catalog. We chose the masked language modeling task with 15% masking rate, also included a binarized orders prediction task. We interleaved the training between these two tasks, with each batch consisting of only one task. We utilized one year of Walmart search behavioral data for the binarized orders prediction training.
@@
Cross Encoder Architecture: We finetune our cross-encoder model using human-labeled query-item relevance pairs and aim to classify the relevance of these pairs. As shown in Figure (1), we begin by merging all query-item features into a single concatenated string. This string is then fed into the BERT encoder. After the BERT encoder pass, we apply dropout and then non-linear layers to transform the embedding into a three-dimensional classification vector. Lastly, we apply a Softmax layer to convert this vector into a three-class relevance probability vector, which is optimized using the cross-entropy loss.
3.2.2. DeBERTa
DeBERTa-v3-large is used, similar to the
BERT-base implementation described above.
3.2.3. Llama2
@@
We finetuned Llama2-7B as a part of cross-encoder architecture for predicting relevance between query and items. We experimented with finetuning a classifier (a dense classification layer superimposed on Llama2 7B), with and without the LoRA implementation. With Low Rank Adaptation of Llama we explored a range of rank and $\alpha$s that dictate the size of linear approximation to the model, and the ratio to which the linear differential approximations were included in training (Hu et al., 2022). The classification itself is based on the “last token” that is not a padding token. We ran experiments with increasing the size of model parameters used in training, based on LoRA rank, to find the optimal model parameter size for classification based on number of labels and training data points.
3.2.4. Mistral
@@
Similar to BERT and Llama, we tried full fine tuning as well as LoRA fine tuning of Mistral-7B-v0.1 using various hyper parameters and learning rate schedulers. We experimented on the following aspects: various learning rate schedulers, including and excluding product description during training and testing phases, product feature dropouts, introducing random negatives, domain adaptation pretraining of Mistral-7B as causal LM on walmart data followed by full fine tuning, and domain adaptation pretraining of LoRA adapters on Mistral-7B as causal LM on Walmart data, followed by LoRA fine tuning of the domain adapted model.
@@ 4. Results: Training Findings, Discussion
4.1. BERT
Table 8 contains the results for BERT-base and DeBERTa-v3-large. Models B1., B2. and B3. in Table 8 differ only in drop out rates from $0.1$ to $0.05$, with little change in outcome. Model B4. is trained starting from a Walmart pretrained BERT, with no class weights (unlike models B1. to B3.) and uses variation # 1 in table 5 with no verbose item description used in training. Models B5. and B6., however, do use verbose description (option #3 in Table 8 ) and as a result show a boost in performance compared to models B3. and B4. Finally models B7. to B11. use public version of BERT, for finetuning, instead of pretrained ones, and show degradation in performance compared to models B1. to B6.
DeBERTa experiments D1. and D2. present LoRA trainings with rank=64, $\alpha$=16, on all linear modules and a learning rate of $41e-$ 4 and learning rate scheduler with linear warmup. The comparison shows that including item descriptions has improved the model’s performance.
4.2. Llama
Table 6 presents the results of experiments involving Llama2 7B. We examined LoRA ranks in the range of 16 to 512, noting that increasing LoRA rank beyond the range of 256 did not improve the model performance. For rank =256, LoRA $\alpha$s closer to half of rank yielded the highest f1 scores. Similar experiments, comparing LoRA dropouts 0.1 and 0.05, found the latter to be preferable. The learning rate for all the Llama experiments is $41e-$ 4, and decreasing the learning rate to $51e-$ 5 did not improve the performance. Compared to the experiment L.3, decreasing $\alpha$ further to 64, worsened, and including class weights, improved the results. Removing added noise to the queries degraded the performance, as well as using item data without description (item data #1 in Table 5). Comparing L4. to L3., we find that resolving the disagreement between two sets of human evaluators via a third vote significantly improves the results on the reformed test set, implying the potential for labelling improvements over the existing human judgment labels. [^5] The procedure for improving the test set label accuracy is described in section 3.1.
Finally, LoRA training shows better performance compared to full finetuning of Llama2 7B with our dataset (L11. versus L4.)
4.3. Mistral
Results are included in Table 7. Including item description during the training helped improve the test metrics, furthermore a comparison between M1. and M2. shows that training on the dataset with higher number of queries per item yields better performance. A similar comparison between Mistral experiments M3.1 and M3.2 shows training on item attribute dataset #2 in Table 5 improves over dataset #3.
@@ 4.3. Mistral
The Domain adaptation pretraining of LoRA adapters, followed by LoRA finetuning did not show a significant improvement over the standalone LoRA finetuning. Also Similar to Llama, LoRA trainings on Mistral result in better performance compared to full model finetuning.
@@
baseline, Bert-base Cross-encoder 0.737 0.695 0.913 0.843
- L0. baseline, Llama2 7B pretrained 0.725 0.662 0.923 0.853
@@
+ L11 Llama2-7B Full finetune (attention_dropout=0.4) 0.786 0.758 0.944 0.889
@@
Table 6. Ablation Studies for Llama2
@@
baseline, Bert-base Cross-encoder 0.737 0.695 0.913 0.843
- M0. baseline, Mistral-7B pretrained 0.669 0.679 0.924 0.853
+ M1. Mistral-7B-v0.1 (attention_dropout=0.4) , #2 item descr dataset 0.789 0.766 0.944 0.891
+ M2. Mistral-7B-v0.1 (attention_dropout=0.4), #3 item descr dataset 0.787 0.752 0.941 0.886
+ M3. Mistral-7B-v0.1 LoRA (r=256, $\alpha$=128, d=0.05, all linear layers), #3 item dataset 0.767 0.739 0.937 0.879
+ M3.1 Mistral-7B-v0.1 LoRA (M3.+ $\alpha$=256+ embedding layer), #2 item dataset 0.800 0.767 0.947 0.895
+ M3.2. Mistral-7B-v0.1 LoRA (same as M3.), #2 item dataset 0.799 0.771 0.947 0.896
+ M3.3. Mistral-7B-v0.1 LoRA (same as M3.), #3 item dataset 0.795 0.768 0.946 0.894
+ M3.4. Mistral-7B-v0.1 LoRA (same as M3.), #2 item dataset, w feature dropout+RN=0.01 0.804 0.770 0.947 0.896
+ M3.4.1. Mistral-7B-v0.1 LoRA (same as M3.), #1 item dataset, w feature dropout+RN=0.01 0.798 0.767 0.945 0.894
+ M3.5. Mistral-7B-v0.1 LoRA (same as M3.),#2 item dataset, w feature dropout+RN=0.3 0.798 0.771 0.947 0.894
+ M3.6. Mistral-7B-v0.1 LoRA (same as M3.), Pretrain 20% + FT prams same as 3.5. 0.786 0.758 0.944 0.889
@@
Table 7. Ablation Studies for Mistral
@@
baseline, Bert-base Cross-encoder 0.737 0.695 0.913 0.843
+ B1. Walmart Pretrained Bert-base with class-weights, dropout=0.1 0.763 0.718 0.921 0.858
+ B2. Walmart Pretrained Bert-base with class-weights, dropout=0.2 0.762 0.719 0.922 0.858
+ B3. Walmart Pretrained Bert-base with class-weights, dropout=0.05 0.763 0.719 0.921 0.858
+B4. Walmart Pretrained Bert-base no class-weights, dropout=0.1 0.763 0.724 0.932 0.873
+B5. Walmart Pretrained Bert-base with class-weights, dropout=0.1, w. item description #2 0.761 0.723 0.925 0.861
+ B6. Walmart Pretrained Bert-base no class-weights, dropout=0.1, w. item description #2 0.765 0.720 0.933 0.873
+ B7. Bert-base no class-weights, dropout=0.1 0.720 0.692 0.922 0.855
+ B8. Bert-base with class-weights, dropout=0.1 0.714 0.691 0.922 0.853
+ B9. Bert-base no class-weights, dropout=0.1 w. item description #3 0.727 0.706 0.927 0.861
+ B10. Bert-base with class-weights, dropout=0.1 w. item description #3 0.724 0.670 0.905 0.831
+11. Bert-base no class-weights, dropout=0.1 w. item description #2 0.725 0.704 0.927 0.860
+ D1. deberta-v3-large training w no item description, LoRA 0.774 0.7380 0.935 0.878
+ D2. deberta-v3-large training w item description # 2, LoRA 0.782 0.753 0.940 0.885
@@
Table 8. Ablation Studies for BERT/DeBERTa, used as Baseline for Larger Models
@@ 6.2. Model Specific Mispredictions
The models consistently showed similar top mispredicted segments in terms of Product-types, Specificity, and Grammar. Mistral and LLama2 performed closely, with Mistral slightly outperforming LLama2. DeBERTa showed a performance similar to the two other LLMs when the search query included Product-type intent, but performed worse in other cases. We sampled and studied unique mispredictions from each of the models and have made the following observations: in all segments, the performance of the Walmart pretrained Bert-Base model consistently lagged behind the other models. While all models made some amount of trivial mispredictions, Mistral and Llama2 encountered this in cases where the queries were not correctly worded or when dealing with numerical ranges. However, DeBERTa and BERT-base models also showed this issue when attributes were not explicitly stated in the product title or queries. In a few cases, DeBERTa and BERT-base models made mispredictions when the query was verbose while Mistral and Llama did not.

### v2

@@ 1. Introduction
Our results in section (3) show that Low Rank Adaptation (LoRA) (Hu et al., 2022) of intermediate-sized with the training datasets sized in the order of $10^{6}$ perform better than full model finetuning. This observation emphasizes the tradeoff between the model and data size. For smaller training datasets, LoRA is expected to align with data more efficiently than full finetuning. We also find that including additional item textual data in the form of description, as expected, improves performance in larger multi-billion parameter models, but not in smaller ones of size in the order of hundred million parameters.
Our training dataset is comprised of $6,055,251$ QIPs, $708,895$ queries and $2,664,741$ items. The dataset, similar to those in the public domain (e.g. (Reddy et al., 2022)), is imbalanced towards the most relevant label, but unlike the ESCI dataset introduced in (Reddy et al., 2022)) has an almost complete coverage of longer item description information. [^1] The dataset introduced in (Reddy et al., 2022) has an “item_description” coverage close to 48%. Our relevance models are classifiers that are trained on inputs of concatenated item textual features and query along with relevance labels. On a three-class graduation of relevance labels (relevant, related, irrelevant) we achieve micro f1 scores of 88.8% and 89.6% with two larger language models. There are improvements up to 4.6% in f1 micro scores compared to the performance of pretrained off-the-shelf decoder-only LLMs (see tables 6, 7, and 8).
@@
We also demonstrate the regions of LoRA rank and $\alpha$, as well as LoRA dropouts, where two 7B LLMs perform best, showing that, unlike conventional wisdom, with the size of the training dataset and LLMs in question, $\alpha$s in the range of 1 to 2 times LoRA rank perform better than alternatives. LoRA finetuning, in turn, results in higher f1 scores than full model finetuning in our trainings.
@@ 3.2. Relevance Models
We have employed 7B versions of two open-source Large Language Modles henceforth called LLM3 (Touvron et al., 2023) and LLM4 (Jiang et al., 2023) in our training along with bidirectional encoder transformer model 1 (Devlin et al., 2018),(Abolghasemi et al., 2022), (Reimers and Gurevych, 2019), henceforth called LLM1, and bidirectional encoder transformer model 2 (He et al., 2023a) henceforth called LLM1, as baselines for larger language models.
@@
In the following, we present the model training procedure for a number of LLMs discussed.
3.2.1. LLM1
@@
Previous studies have shown that the cross encoder based LLM1 model (Luan et al., 2021) performs well on solving semantic retrieval problems. To address the specific challenge of e-commerce product relevance search, we have employed a cross encoder model in our design. We experimented with finetuning our model using two variations of the LLM1 model as the starting checkpoints: 1) the LLM1-Base uncased model from Huggingface, and 2) this same model further pretrained on the Walmart Search corpus.
@@
Walmart Pretrained LLM1: Since the LLM1-base model was trained on general domain text data, we aimed to enhance its performance on e-commerce related texts by conducting additional language model training using Walmart product catalog. We chose the masked language modeling task with 15% masking rate, also included a binarized orders prediction task. We interleaved the training between these two tasks, with each batch consisting of only one task. We utilized one year of Walmart search behavioral data for the binarized orders prediction training.
@@
Cross Encoder Architecture: We finetune our cross-encoder model using human-labeled query-item relevance pairs and aim to classify the relevance of these pairs. As shown in Figure (1), we begin by merging all query-item features into a single concatenated string. This string is then fed into the LLM1 encoder. After the LLM1 encoder pass, we apply dropout and then non-linear layers to transform the embedding into a three-dimensional classification vector. Lastly, we apply a Softmax layer to convert this vector into a three-class relevance probability vector, which is optimized using the cross-entropy loss.
3.2.2. LLM2
LLM2, an improvement upon LLM1, is used, similar to the LLM1 implementation described above.
3.2.3. LLM3
@@
We finetuned LLM3-7B as a part of cross-encoder architecture for predicting relevance between query and items. We experimented with finetuning a classifier (a dense classification layer superimposed on LLM3 7B), with and without the LoRA implementation. With Low Rank Adaptation of LLM3 we explored a range of rank and $\alpha$s that dictate the size of linear approximation to the model, and the ratio to which the linear differential approximations were included in training (Hu et al., 2022). The classification itself is based on the “last token” that is not a padding token. We ran experiments with increasing the size of model parameters used in training, based on LoRA rank, to find the optimal model parameter size for classification based on number of labels and training data points.
3.2.4. LLM4
@@
Similar to LLM1 and LLM3, we tried full fine tuning as well as LoRA fine tuning of LLM4-7B-v0.1 using various hyper parameters and learning rate schedulers. We experimented on the following aspects: various learning rate schedulers, including and excluding product description during training and testing phases, product feature dropouts, introducing random negatives, domain adaptation pretraining of LLM4-7B as causal LM on walmart data followed by full fine tuning, and domain adaptation pretraining of LoRA adapters on LLM4-7B as causal LM on Walmart data, followed by LoRA fine tuning of the domain adapted model.
@@ 4. Results: Training Findings, Discussion
4.1. LLM1
Table 8 contains the results for LLM1 and LLM2. Models B1., B2. and B3. in Table 8 differ only in drop out rates from $0.1$ to $0.05$, with little change in outcome. Model B4. is trained starting from a Walmart pretrained LLM1, with no class weights (unlike models B1. to B3.) and uses variation # 1 in table 5 with no verbose item description used in training. Models B5. and B6., however, do use verbose description (option #3 in Table 8 ) and as a result show a boost in performance compared to models B3. and B4. Finally models B7. to B11. use public version of LLM1, for finetuning, instead of pretrained ones, and show degradation in performance compared to models B1. to B6.
LLM2 experiments D1. and D2. present LoRA trainings with rank=64, $\alpha$=16, on all linear modules and a learning rate of $41e-$ 4 and learning rate scheduler with linear warmup. The comparison shows that including item descriptions has improved the model’s performance.
4.2. LLM3
Table 6 presents the results of experiments involving LLM3 7B. We examined LoRA ranks in the range of 16 to 512, noting that increasing LoRA rank beyond the range of 256 did not improve the model performance. For rank =256, LoRA $\alpha$s closer to half of rank yielded the highest f1 scores. Similar experiments, comparing LoRA dropouts 0.1 and 0.05, found the latter to be preferable. The learning rate for all the LLM3 experiments is $41e-$ 4, and decreasing the learning rate to $51e-$ 5 did not improve the performance. Compared to the experiment L.3, decreasing $\alpha$ further to 64, worsened, and including class weights, improved the results. Removing added noise to the queries degraded the performance, as well as using item data without description (item data #1 in Table 5). Comparing L4. to L3., we find that resolving the disagreement between two sets of human evaluators via a third vote significantly improves the results on the reformed test set, implying the potential for labelling improvements over the existing human judgment labels. [^5] The procedure for improving the test set label accuracy is described in section 3.1.
Finally, LoRA training shows better performance compared to full finetuning of LLM3 7B with our dataset (L11. versus L4.)
4.3. LLM4
Results are included in Table 7. Including item description during the training helped improve the test metrics, furthermore a comparison between M1. and M2. shows that training on the dataset with higher number of queries per item yields better performance. A similar comparison between LLM4 experiments M3.1 and M3.2 shows training on item attribute dataset #2 in Table 5 improves over dataset #3.
@@ 4.3. LLM4
The Domain adaptation pretraining of LoRA adapters, followed by LoRA finetuning did not show a significant improvement over the standalone LoRA finetuning. Also Similar to LLM3, LoRA trainings on LLM4 result in better performance compared to full model finetuning.
@@
baseline, LLM1-base Cross-encoder 0.737 0.695 0.913 0.843
- L0. baseline, LLM3 7B pretrained 0.725 0.662 0.923 0.853
@@
+ L11 LLM3-7B Full finetune (attention_dropout=0.4) 0.786 0.758 0.944 0.889
@@
Table 6. Ablation Studies for LLM3
@@
baseline, LLM1-base Cross-encoder 0.737 0.695 0.913 0.843
- M0. baseline, LLM4-7B pretrained 0.669 0.679 0.924 0.853
+ M1. LLM4-7B-v0.1 (attention_dropout=0.4) , #2 item descr dataset 0.789 0.766 0.944 0.891
+ M2. LLM4-7B-v0.1 (attention_dropout=0.4), #3 item descr dataset 0.787 0.752 0.941 0.886
+ M3. LLM4-7B-v0.1 LoRA (r=256, $\alpha$=128, d=0.05, all linear layers), #3 item dataset 0.767 0.739 0.937 0.879
+ M3.1 LLM4-7B-v0.1 LoRA (M3.+ $\alpha$=256+ embedding layer), #2 item dataset 0.800 0.767 0.947 0.895
+ M3.2. LLM4-7B-v0.1 LoRA (same as M3.), #2 item dataset 0.799 0.771 0.947 0.896
+ M3.3. LLM4-7B-v0.1 LoRA (same as M3.), #3 item dataset 0.795 0.768 0.946 0.894
+ M3.4. LLM4-7B-v0.1 LoRA (same as M3.), #2 item dataset, w feature dropout+RN=0.01 0.804 0.770 0.947 0.896
+ M3.4.1. LLM4-7B-v0.1 LoRA (same as M3.), #1 item dataset, w feature dropout+RN=0.01 0.798 0.767 0.945 0.894
+ M3.5. LLM4-7B-v0.1 LoRA (same as M3.),#2 item dataset, w feature dropout+RN=0.3 0.798 0.771 0.947 0.894
+ M3.6. LLM4-7B-v0.1 LoRA (same as M3.), Pretrain 20% + FT prams same as 3.5. 0.786 0.758 0.944 0.889
@@
Table 7. Ablation Studies for LLM4
@@
baseline, LLM1-base Cross-encoder 0.737 0.695 0.913 0.843
+ B1. Walmart Pretrained LLM1-base with class-weights, dropout=0.1 0.763 0.718 0.921 0.858
+ B2. Walmart Pretrained LLM1-base with class-weights, dropout=0.2 0.762 0.719 0.922 0.858
+ B3. Walmart Pretrained LLM1-base with class-weights, dropout=0.05 0.763 0.719 0.921 0.858
+B4. Walmart Pretrained LLM1-base no class-weights, dropout=0.1 0.763 0.724 0.932 0.873
+B5. Walmart Pretrained LLM1-base with class-weights, dropout=0.1, w. item description #2 0.761 0.723 0.925 0.861
+ B6. Walmart Pretrained LLM1-base no class-weights, dropout=0.1, w. item description #2 0.765 0.720 0.933 0.873
+ B7. LLM1-base no class-weights, dropout=0.1 0.720 0.692 0.922 0.855
+ B8. LLM1-base with class-weights, dropout=0.1 0.714 0.691 0.922 0.853
+ B9. LLM1-base no class-weights, dropout=0.1 w. item description #3 0.727 0.706 0.927 0.861
+ B10. LLM1-base with class-weights, dropout=0.1 w. item description #3 0.724 0.670 0.905 0.831
+11. LLM1-base no class-weights, dropout=0.1 w. item description #2 0.725 0.704 0.927 0.860
+ D1. LLM2 training w no item description, LoRA 0.774 0.7380 0.935 0.878
+ D2. LLM2 training w item description # 2, LoRA 0.782 0.753 0.940 0.885
@@
Table 8. Ablation Studies for LLM1/LLM2, used as Baseline for Larger Models
@@ 6.2. Model Specific Mispredictions
The models consistently showed similar top mispredicted segments in terms of Product-types, Specificity, and Grammar. LLM4 and LLM3 performed closely, with LLM4 slightly outperforming LLM3. LLM2 showed a performance similar to the two other LLMs when the search query included Product-type intent, but performed worse in other cases. We sampled and studied unique mispredictions from each of the models and have made the following observations: in all segments, the performance of the Walmart pretrained LLM1-Base model consistently lagged behind the other models. While all models made some amount of trivial mispredictions, LLM3 and LLM4 encountered this in cases where the queries were not correctly worded or when dealing with numerical ranges. However, LLM2 and LLM1-base models also showed this issue when attributes were not explicitly stated in the product title or queries. In a few cases, LLM2 and LLM1-base models made mispredictions when the query was verbose while LLM3 and LLM4 did not.